OpenStudy (miss.rose):

Can you help me with 5 Mixed Operation Applications questions? I will medal and fan. Please help! @rebeccaxhawaii

rebeccaxhawaii (rebeccaxhawaii):

yes

OpenStudy (miss.rose):

Okay, thanks! First Question: The difference between twice a number and two is 10. Write and solve an equation to determine the number A. x = 4 B. x = 6 C. x= 7 D. x= 9

OpenStudy (anonymous):

Do you have any idea on how to setup the equation?

OpenStudy (miss.rose):

Not really.

OpenStudy (anonymous):

"Difference" implies that you need a "minus". And twice is 2 times something (x). So it would look like this \[2x-2=10\] Does that make sense ?

rebeccaxhawaii (rebeccaxhawaii):

yes you 10-2=8 divide by 2 giving you what ?

OpenStudy (miss.rose):

4?

OpenStudy (anonymous):

Yea, correct. \[x=4\]

rebeccaxhawaii (rebeccaxhawaii):

good job

OpenStudy (miss.rose):

The sum of 1 and 3 times a number is 22. Write and solve an equation to determine the number. A. x=4 B. x=6 C. x=7 D. x=9 Can you walk me through it again?

OpenStudy (anonymous):

Setup the equation. The sum: implies additions, and then we have 3 times a number. So the equation would look like this:\[1+3x=22\]

OpenStudy (miss.rose):

7?

OpenStudy (anonymous):

Correct

OpenStudy (miss.rose):

Okay. Next Question: Machine A makes 50 more than 4 times the number of parts that Machine B makes each day. If Machine A makes 650 parts each day, how many parts does Machine B make each day? A. 150 B. 175 C. 700 D. 2650 I need help again.

OpenStudy (miss.rose):

Someone?

OpenStudy (anonymous):

So you can make the following equations: \[A=50+4B\] \[A=650\] Insert and solve B :)

OpenStudy (miss.rose):

b = 150?

OpenStudy (anonymous):

Yes :)

OpenStudy (miss.rose):

Okay, thanks! Next Question: Matthew is 5 years older than twice the age of his sister. If Matthew is 13, how old is his sister? A. 3 This one or B. 4 This one? C. 9 D. 31

OpenStudy (anonymous):

Do you have any idea how the equation would look?

OpenStudy (miss.rose):

No. Please help. I have to leave soon.

OpenStudy (anonymous):

\[M=5+2S\] \[M=13\]

OpenStudy (miss.rose):

s = 4?